find the x and y intercepts for the following equation:
f(x)=x^2-4x+2

Answer by nyc_function(2741)

f(x)=x^2-4x+2
To find the x-intercept, let y = 0 and solve for x.
NOTE: f(x) = y and vice-versa.
0 = x^2 - 4x + 2
After doing the math using the quadratic formula, I got x = [2 + sqrt{2}] and
x = [2 - sqrt{2}].
To find the y-intercept, let x = 0 and solve for y.
y = (0)^2 - 4(0) + 2
y = 0 - 0 + 2
y = 2
Done!